OpenStudy (anonymous):

A spring is moving up and down according to the function s(t) = -4cos(t). Part a) Find the first derivative and calculate the velocity of the spring at t = 3pi/4 Part b) Is the spring moving upward or downward at this point? Justify your answer.

OpenStudy (anonymous):

do you know the derivative of cosine?

OpenStudy (anonymous):

-sin(x)

OpenStudy (anonymous):

at 3pi/4 what does sine equal?

OpenStudy (anonymous):

yes so in this case the derivative of \(-4\cos(x)\) is \(4\sin(x)\)

OpenStudy (zzr0ck3r):

so the funtion is 4sin(x)|x = (3pi)/4 4sin((3pi/4) = 4(sqrt(2)/2) = 2sqrt(2)

OpenStudy (anonymous):

I hateee these questions why does physics and math have to conflict on so many levels

OpenStudy (zzr0ck3r):

how does in conflict?

OpenStudy (anonymous):

replace \(x\) by \(\frac{3\pi}{4}\) and see what you get that is the velocity if it is positive, going up if it is negative, going down

OpenStudy (anonymous):

not in this case but however if you had a negative velocity

OpenStudy (zzr0ck3r):

just means pointing down

OpenStudy (zzr0ck3r):

negative vector is direction not value

OpenStudy (anonymous):

and math doesn't take that into account however physics does and it bothersme

OpenStudy (zzr0ck3r):

I dont understand what you are saying, a negative vector is fine in physics and mathematics

OpenStudy (anonymous):

Physicists say that velocity cannot be negative it can only be zero or a positive magnitude from zero in a direction, where as mathematicians allow negative velocity, however it means that it is traveling in the oppositve direction

OpenStudy (zzr0ck3r):

no it says speed cant be negative, speed = abs(velocity)

OpenStudy (anonymous):

that's another way of saying it yes

OpenStudy (zzr0ck3r):

speed is magnitude of vilocity

OpenStudy (anonymous):

velociy is speed and it's direction, you cannot have a negative speed

OpenStudy (zzr0ck3r):

but we use negative vilocity all over the place in physics but its arbitrary because you can just flip the whole system in most cases, so sign is an indication of the origin of the axis which you supply

OpenStudy (zzr0ck3r):

right velocity is speed and direction, and direction being negative says nothing about speed being negative.

OpenStudy (zzr0ck3r):

if you hit a ball in the air and its falling and you are still on the ground its velocity is negative in respect to where you are, but if you rise to the vertex of where the ball went, and droped the ball so it took the same path down, the velocity is now positive with respect to where you are.

OpenStudy (queelius):

Velocity is a vector; speed is a scalar. A vector has a direction, a scalar does not. If our problem is in a one-dimensional space, then we only have two possible directions; the sign, therefore, is used to denote the direction.

OpenStudy (anonymous):

what if it is n-dimensional space

OpenStudy (zzr0ck3r):

it still denotes direction with respect to origin

OpenStudy (zzr0ck3r):

just the opposite direction of it being pos

OpenStudy (zzr0ck3r):

just with 2d you can say up down, left, right...

OpenStudy (zzr0ck3r):

-<1,1,1> points in the direction <-1,-1,-1>